# Quantumness and Randomness of Quantum Random Number Generators via Photon Statistics

### v2

@@ Abstract
Randomness is one of the essential components in many fields including cryptography and simulations. Several quantum random number generator (QRNG) models have been proposed to produce random numbers, which, due to the quantum theory, are more secure than their classical counterparts. Many QRNG devices are commercially available as off-the-shelf black-box devices. Computationally, it is not possible to distinguish between a PRNG and a QRNG just by observing their outputs. The common practice for testing quantumness is a direct comparison between the mean and the variance of the experimental photon count. However, statistically, this is not a feasible solution either, because of finite sample size. A QRNG that uses single photons as the quantum source produces true quantum random numbers. Since single photons follow sub-Poissonian statistics, by determining the underlying distribution one can conclude whether a QRNG is truly quantum or not. In this work, we point out the limitations of existing methods of such a decision-making processes, and propose a more efficient two-fold statistical method, which can ensure whether an optical source is quantum or not up to a desired confidence level. Also, QRNGs can not produce true random numbers without deterministic classical post-processing. In this work, we also show that the two models of QRNGs, one producing random numbers from exponential distribution and the other from uniform distribution, become essentially similar under device noise. Detector outputs of both the above models can be tested to quantify the randomness coming from the quantum source of a QRNG, which in turn, dictates how much post-processing is required to produce good random numbers retaining quantumness. In this context, we also derive a relation when the underlying sampling distributions of the QRNGs will be $\epsilon$-random. Depending on this relation, a suitable post-processing algorithm can be chosen.
@@ I Introduction
Random numbers play a crucial role in numerous applications, including computational methods like Monte Carlo simulations and programming Carlson et al. (2015); Metropolis and Ulam (1949), testing principles of physics and chemistry Bell (1964), as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett et al. (1992); Shannon (1949), extending to commercial uses such as slot machines and lottery games Konstantinou et al. (2004); Yamanaka and Nakano (2017). In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the other hand, the randomness of the random numbers generated by quantum random number generators (QRNG) relies on the theory of quantum physics Ma et al. (2016). Different technologies, including shot noise, tunneling, fluctuation, photon arrival time, etc. have been proposed for QRNGs. Many manufacturers are using these technologies to produce their QRNGs. Herrero-Collantes and Garcia-Escartin (2017); ID Quantique (2001); Qutools (2010); QNu Labs (2018); Quintessence Labs (2022); Quside (2023); Quantum CTek (2023); Mars Innovation and Department of Electrical and Computer Engineering & Department of Physics, University of Toronto (2023); Quntum Computing Inc. (2023). These QRNGs are commercially available as off-the-shelf black-box devices. Statistically, it is not feasible to distinguish between a PRNG and a QRNG just by observing their outputs. So a pertinent question arises: how do we test whether an alleged QRNG is truly quantum or not?
Let $S_{k}=\{0,1\}^{k}$ be the set of all $k$-bit binary strings. Then $S_{k}$ contains $2^{k}$ elements. A $n$-bit string is called true (or uniformly) random if and only if for any $k$-bit subset ($k\leq n$) $x$ of this string, the probability that $x=y$ for some $y\in S_{k}$ is $\frac{1}{2^{k}}$. Existing time-bin-based QRNG models produce random numbers from exponential Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010) or uniform Nie et al. (2014) distributions if the underlying optical devices perform ideally. But, in practice, they cannot be ideal. Therefore, all of the QRNGs require some classical post-processing using deterministic functions. These post-processings use a classical source of randomness to increase randomness in the raw data generated by the QRNGs. Therefore, the amount of classical post-processing should be minimal to ensure the quantum source of the randomness in the final random numbers. For this, the raw number (before classical post-processing) generated by QRNG should be as close to uniformly random as possible.
In this article, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest two-fold statistical methods to verify whether the source of the sample is quantum or not. Here we assume that the sample only contains the counts detected by the photon detector without any kind of processing. This sample may also contain the dark count of the detector.
@@
Alléaume et al. Alléaume et al. (2004) discussed photon statistics of single-photon sources using the Mandel parameter given by (Mandel, 1979, Equation (11a)). They calculated values of the Mandel parameter from observed photon count timestamps. They concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources. Some recent experimental works Jones et al. (2023); Peřina et al. (2024a, b); Bush and Børkje (2024) also made the same conclusion. Although analytically it is true, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same as the mean and the variance of the underlying distribution. Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
@@
Another approach of testing quantumness of light is to experimentally observe photon antibunching Paul (1982). However, although a stream of antibunched photons indicate the non-classicality nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons Zou and Mandel (1990); Fox (2006). We explain this in more detail in Appendix A.
Our main contribution in this article is to devise a two-fold statistical method to overcome the limitations of the above two existing approaches. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level. This can validate the quantumness of light more accurately and efficiently.
@@
As another contribution, we derive the photon statistics for the photon arrival-time-based QRNGs in practical scenarios. We use these statistics to test the randomness of the generated random number. Heavy post-processing should be applied to produce truly random numbers if these raw numbers are less random. Therefore, photon statistics for QRNGs, which we discuss in this article, can be considered as one measure of randomness from quantum sources for photon arrival-time-based QRNGs. Depending on this measure, a suitable post-processing algorithm can be chosen to retain the maximum amount of randomness from the quantum source. For the practical scenario, we consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
II Classical vs Quantum Light: Photon-counting Statistics
@@
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols. In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, the photon-counting statistics have an important role in determining whether the light falls in the quantum regime or not.
@@
Classical wave theory of light follows Poissonian and super-Poissonian statistics Fox (2006). On the other hand, the particle nature of light, considered as a sequence of single photons follows sub-Poissonian distributions. Therefore, the sub-Poissonian nature of a photon source is a clear indication of its quantumness Davidovich (1996); Paraïso et al. (2021).
@@
where $\mu$ is the mean number of photons per unit time. One can show that for this distribution, the variance of the number of photons per unit time is also $\mu$.

### v3

# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

@@ Abstract
Despite the availability of commercial QRNG devices, distinguishing between PRNG and QRNG outputs computationally remains challenging. This paper presents two significant contributions from the perspectives of QRNG manufacturers and users. For manufacturers, the conventional method of assessing the quantumness of single-photon-based QRNGs through mean and variance comparisons of photon counts is statistically unreliable due to finite sample sizes. Given the sub-Poissonian statistics of single photons, confirming the underlying distribution is crucial for validating a QRNG’s quantumness. We propose a more efficient two-fold statistical approach to ensure the quantumness of optical sources with the desired confidence level. Additionally, we demonstrate that the output of QRNGs from exponential and uniform distributions exhibit similarity under device noise, deriving corresponding photon statistics and conditions for $\epsilon$-randomness.
From the user’s perspective, the fundamental parameters of a QRNG are quantumness (security), efficiency (randomness and random number generation rate), and cost. Our analysis reveals that these parameters depend on three factors, expected photon count per unit time, external reference cycle duration, and detection efficiency. A lower expected photon count enhances security but increases cost and decreases the generation rate. A shorter external reference cycle boosts security but must exceed a minimum threshold to minimize timing errors, with minor impacts on cost and rate. Lower detection efficiency enhances security and lowers cost but reduces the generation rate. Finally, to validate our results, we perform statistical tests like NIST, Dieharder, AIS-31, ENT etc. over the data simulated with different values of the above parameters. Our findings can empower manufacturers to customize QRNGs to meet user needs effectively.
@@ I Introduction
Random numbers have important roles in numerous applications, including computational methods like Monte Carlo simulations and programming Carlson et al. (2015); Metropolis and Ulam (1949), testing principles of physics and chemistry Bell (1964), as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett et al. (1992); Shannon (1949), extending to commercial uses such as slot machines and lottery games Konstantinou et al. (2004); Yamanaka and Nakano (2017). In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the contrary, the randomness of a true random number generator (TRNG) is rooted in some physical entropy source Schindler (2009); Stipčević and Koç (2014). A quantum random number generator (QRNG) uses the quantum particles and the principles of quantum physics Ma et al. (2016) to produce random numbers.
Statistically, it is not feasible to distinguish between a PRNG and a QRNG just by analyzing their outputs. Statistical tests like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), AIS-$31$ Killmann and Schindler (2011), ENT Walker (2008) etc. are available to test the randomness of a random number generator (RNG). However, from the results of these tests, it is not possible to say whether an RNG is quantum or not. So a pertinent question arises:
Q1: How do we test whether an alleged QRNG is truly quantum or not?
Different technologies, including photon arrival time Stipčević and Rogina (2007); Dynes et al. (2008); Wayne and Kwiat (2010); Michael A. Wayne and Kwiat (2009); Wahl et al. (2011); Nie et al. (2014); Khanmohammadi et al. (2015); Zhang et al. (2021), tunneling Bernardo-Gavito et al. (2017); Aungskunsiri et al. (2021), fluctuation Gabriel et al. (2010); Haw et al. (2015); Guo et al. (2019); Zhou et al. (2019); Zheng et al. (2019); Gehring et al. (2021), phase noise Qi et al. (2010); Guo et al. (2010); Shen et al. (2010); Xu et al. (2012); Nie et al. (2015); Yang et al. (2016); Huang et al. (2020), etc. have been proposed to produce QRNGs Herrero-Collantes and Garcia-Escartin (2017); Mannalatha et al. (2023). Several manufacturers such as ID Quantique ID Quantique (2001), Qutools Qutools (2010), QNuLabs QNu Labs (2018), QuintessenceLabs Quintessence Labs (2022), Mars Innovation Mars Innovation and Department of Electrical and Computer Engineering & Department of Physics, University of Toronto (2023), Quantum Computing Inc. Quntum Computing Inc. (2023), Quside Quside (2023), QuantumCTek Quantum CTek (2023) are producing QRNGs following these technologies. These QRNGs are commercially available as off-the-shelf black-box devices. As the photon arrival-time-based QRNGs are the most general and most of the companies follow this approach, in this work we try to answer the Q1 with respect to this type of QRNGs.
@@
In most of the analyses Jones et al. (2023); Peřina et al. (2024a, b); Bush and Børkje (2024) of the photon arrival-time-based QRNGs, the photon statistics have been discussed either by comparing the mean and variance of the sample distribution or using the negativity of the Mandel parameter Mandel (1979). For example, Alléaume et al. Alléaume et al. (2004) discussed photon statistics of photon sources using the Mandel parameter given by (Mandel, 1979, Eq. (11a)). They calculated values of the Mandel parameter from observed photon count timestamps and concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources. Although analytically it is true, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same as the mean and the variance of the underlying distribution. Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
@@
Another approach to test the quantumness of light is to experimentally observe photon antibunching Paul (1982). However, although a stream of antibunched photons indicates the non-classical nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons Zou and Mandel (1990); Fox (2006). We explain this in more detail in Appendix A. This approach also consists of an equality criteria similar to above photon statistics case. Therefore this is also not enough to decide photon antibunching.
On the other hand, a single-photon source ensures the quantumness of a QRNG Paul (1982); Fox (2006). However, generating single photons are experimentally challenging taskImamoḡlu and Yamamoto (1994); Kim et al. (1999). Various experimental methods Brunel et al. (1999); Lounis and Moerner (2000); Santori et al. (2001, 2002); Vučković et al. (2003); Kuhn et al. (2002); Adam and Mechler (2024); Wang and Wang (2024) have been explored to produce single photons. Existing time-bin-based QRNG models utilize coherent photon sources in conjunction with attenuators to produce random numbers from exponential Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Wahl et al. (2011) or uniform Stipčević and Rogina (2007) distributions if the underlying optical devices perform ideally and the source generates perfectly single photon. But, in practice, this is not the case. Therefore, it is important to explore the photon statistics in practical scenarios.
Also, all of the QRNGs require some classical post-processing using deterministic functions. These post-processings use a classical source of randomness to increase randomness in the raw data generated by the QRNGs. Side-channel attacks Kuznetsov et al. (2019); Park et al. (2019); Chowdhury et al. (2021) can also be performed in this post-processing step to gather information about the random number. Therefore, the amount of classical post-processing should be minimal to ensure the randomness from the quantum source in the final outputs and reduce the possibility of side-channel attacks on classical post-processing. To achieve this, the raw output number (before classical post-processing) generated by QRNG should be as close to uniform random as possible. As all the practical devices are prone to error, this raises an important question:
Q2: How to increase the randomness from the quantum source in the presence of device errors?
Besides quantumness and randomness, the random number generation rate and the cost of the QRNG are also important parameters from the user’s perspective. The trade-off between the randomness from the quantum source, the random number generation rate and the cost of the QRNG depends on the users’ requirements. For example, in the case of secure communication, the randomness from the quantum source should have the highest priority, and also the generation rate should be high. On the other hand, in token generation or lottery games, a low rate of random number generation would be sufficient. Therefore, to fulfill the user’s requirement, the answers to the following questions are required.
Q3: How to choose the optical parameters (a) the photon numbers in unit time, (b) the time cycle and (c) the detection efficiency of a QRNG depending on the user’s requirements?
In this article, we aim to evaluate the quantumness of photon arrival-time-based QRNG from the manufacturer’s perspective as answers to Q1 and Q2 and randomness, random number generation rate, and cost from the user’s perspective as the answer to Q3. For the manufacturer, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest a two-fold statistical method to verify whether the source of the sample is quantum or not, overcoming the limitations of the existing approaches discussed above. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level.
@@
As another contribution, also from the manufacturer’s perspective, we derive the photon statistics for two-photon arrival-time-based QRNGs in practical scenarios. These photon statistics for QRNGs can be considered as a measure of randomness from quantum sources for photon arrival-time-based QRNGs. Depending on this measure, a suitable post-processing algorithm can be chosen to retain the maximum amount of randomness from the quantum source. For the practical scenario, we consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
From the user’s perspective, we see that the maximum randomness from the quantum source may not be achieved due to some practical constraints related to the random number generation rate, the timing error in registering the photon detection time and the cost of the QRNG. In particular, we show the following.
  • If the expected number of photons within a unit of time is small, the randomness from the quantum source would be high. However, it would increase the cost of the QRNG, and also, it would reduce the random number generation rate.
  • Although a small external reference cycle of the QRNG increases the randomness from the quantum source, to minimize the timing error, it should be higher than a specific value. This reference cycle does not have much effect on the cost or random number generation rate of the QRNG.
  • A detector with less efficiency increases the randomness from the quantum source and also reduces the cost of the QRNG. However, it also reduces the random number generation rate.
Finally, we simulated random numbers depending on photon statistics we have derived, and performed statistical tests like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), AIS-$31$ Killmann and Schindler (2011) and ENT Walker (2008) on those random numbers. We show that these statistical tests support our theoretical deduction regarding randomness from quantum sources.
This paper is structured as follows. In Section II we perform some preliminary discussion about photon statistics, some statistical tools used in this work and common randomness testing suites. In Section III and IV we revisit single-photon-based QRNGs and coherent pulse-based QRNGs respectively and mention their weakness in present days. Section V reveals our two-fold proposal to check the quantumness of a quantum device as an answer to Q1. Detailed discussion on coherent pulse-based QRNGs has been performed in Section VI to answer Q2. Section VII answers Q3 by separately analyzing the effects of the parameters mentioned in that question. Results from different randomness testing suites for our simulated random numbers based on the analysis of Section VI are given in Section VIII. Finally, in Section IX we conclude our work.
II Preliminaries
In this section we first discuss photon statistics. We also discuss how these photon statistics distinguish quantum light from classical light. After that, we mention some statistical tools used in this work. Then we briefly discuss common randomness testing suites for testing the randomness of random numbers
II.1 Photon Statistics
@@
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols Bennett et al. (1992); Shannon (1949). In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, the photon-counting statistics have an important role in determining whether the light falls in the quantum regime or not.
@@
Classical wave theory of light follows Poissonian and super-Poissonian statistics Fox (2006). On the other hand, the sub-Poissonian nature of a photon source is a clear indication of its quantumness Davidovich (1996); Paraïso et al. (2021).
@@
where $\mu$ is the mean number of photons per unit time. It can be shown that for this distribution, the variance of the number of photons per unit time is also $\mu$.
II.2 Statistical Tools
II.2.1 Interval Estimation
In statistics, an interval $(a,b)$ is called the confidence interval for a distribution parameter $\Theta$, if $P(A<\Theta<B)=1-\epsilon$ for some predefined $\epsilon\in(0,1)$, where $A,B$ are the random variables corresponding to the statistics $a,b$ (they are functions of sample values) respectively. The quantity $1-\epsilon$ is called confidence level of the interval. The procedure of finding the confidence interval for the parameter $\Theta$ is called interval estimation. A statistic $a$ is called an unbiased estimate of an unknown parameter $\Theta$ if the expectation $\langle A\rangle=\Theta$. A statistic $a$ is called a consistent estimate of a parameter $\Theta$ if for every real $\epsilon>0$, $\lim\limits_{n\to\infty}P(|A-\Theta|\geq\epsilon)=0$, where $n$ is the sample size.
II.2.2 Hypothesis Testing
In statistics, a hypothesis is a statement or an assumption about a population distribution. this may be right or wrong. Hypothesis testing is a procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample Casella and Berger (2002). A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
II.2.3 Goodness-of-fit Test
Goodness-of-fit test is a statistical procedure to decide whether sample data comes from a particular distribution or not. This is a specific hypothesis testing where
$H_{0}$: Sample data comes from a distribution to be tested.
$H_{a}$: Sample data not from that distribution.
Chi-square ($\chi^{2}$) testCochran (1952), Kolmogorov–Smirnov test (KS test) Massey (1951), and Anderson–Darling test Anderson and Darling (1952) are some examples of the goodness-of-fit test Dodge (2008a). Out of these tests, the last two tests are designed for continuous distributions only. However, the KS test is modified later to consider discrete distributions as well Conover (1972). The $\chi^{2}$ test has been found more robust than the KS test for the continuous distributions Slakter (1965).
II.3 Randomness Testing Suites
Randomness testing is very important to ensure whether an RNG is producing perfect random numbers or not. It tells us about the uniformity of the underlying distribution that the outputs of an RNG follow.
Definition 1.
Let $S_{k}=\{0,1\}^{k}$ be the set of all $k$-bit binary strings containing $2^{k}$ elements. A $n$-bit string, $X=x_{1}x_{2}\dots x_{n},\ x_{i}\in\{0,1\}$ for all $i$, is called perfect (or uniformly) random if and only if for any $k$-bit ($k\leq n$) sub-string $x=x_{j}x_{j+1}\dots x_{j+k}$, for some $j\leq n-k$, of this string, the probability that $x=y$ for any $y\in S_{k}$ is $\frac{1}{2^{k}}$. That is,
$X=x_{1}x_{2}\dots x_{n},x_{i}\in\{0,1\}\ \forall i,\text{ is true random}$
$\iff$ $\forall k<n,\forall j\leq n-k\text{ and }\forall y\in S_{k},\;P(x=y)=\frac{1}{2^{k}},$
where $x=x_{j}x_{j+1}\dots x_{j+k}$.
This definition of a perfect random number uses probability over all possible combinations of the bit string to ensure randomness. However, in a practical case, when we have only a finite sample, this probability cannot be perfectly calculated. Hence this definition cannot be directly used to test the randomness of random numbers.
To overcome this limitation, several randomness testing suites like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), AIS-$31$ Killmann and Schindler (2011), ENT Walker (2008) etc. have been proposed. Recently, Foreman et al. also suggested some test suite Foreman et al. (2024) for testing randomness. Each of these suits consists of multiple tests to decide the randomness of the input data. For each test, they calculate the value of the test statistic, and depending on these values the randomness is decided. NIST and Dieharder perform hypothesis testing with the null hypothesis that the number is random, AIS-31 performs interval estimation, and ENT directly returns the test statistic values. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation Metropolis and Ulam (1949) to validate the randomness.
III Revisiting QRNG based on Single-photon Source
QRNG, based on single photons, uses $/50$ 50 beam splitter Jennewein et al. (2000), where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Fig. 1.
[…]
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section V. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
Each bit of the random number generated with this single-photon-based QRNG has probability $\frac{1}{2}$ of being 00 or $1$. So, ideally it produces uniform random numbers. However, any defect in the beam-splitter or the detectors may change this probability. For example, consider an $R/T$ beam-splitter and the detectors $D_{i}$ with detection efficiency $d_{i}\in(0,1]$. Suppose, $D_{0}$ is placed on the path of the transmitted light, and $D_{1}$ is placed on the path of the reflected light. Then the probability of 00 in the generated random number, that is, the probability of detecting a photon by detector $D_{0}$ would be
$P(D_{0})=\frac{d_{0}}{d_{0}+d_{1}}\frac{T}{100}.$ (5)
Similarly, the probability of $1$ would be
$P(D_{1})=\frac{d_{1}}{d_{0}+d_{1}}\frac{R}{100}.$ (6)
If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST, Dieharder, AIS-$31$, ENT etc. may be used. See Section VIII for a discussion on these.
IV Revisiting QRNG based on Coherent Pulse
It is well known that producing single photon states is experimentally challenging Imamoḡlu and Yamamoto (1994); Kim et al. (1999). However, different experimental approaches Brunel et al. (1999); Lounis and Moerner (2000); Santori et al. (2001, 2002); Vučković et al. (2003); Kuhn et al. (2002); Adam and Mechler (2024); Wang and Wang (2024) have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval Argillander et al. (2023). In this process, the photons follow Poissonian statistics. A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Fig. 2.
SourceAttenuatorSPDFPGARandom Numbers
Figure 2: A basic block diagram of the photon arrival time-based QRNG using weak coherent pulses. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval becomes negligible Argillander et al. (2023). A SPD detects these photons and the detection time (photon arrival time) has been registered. The Field Programmable Gate Array (FPGA) provides random numbers based on the registered time stamps.
If $\lambda$ is the wavelength of the photon wave and $n$ photons are present within $T$ time, the corresponding photon power would be
$P=\frac{n\hbar c}{\lambda T}=\frac{\mu\hbar c}{\lambda},$ (7)
where $\hbar$ is the Planck’s constant and $c$ is the speed of light in vacuum. Let $P_{a}$ be the photon power coming out from the attenuators. Then, the expected number of photons per unit time is given by
$\mu=\frac{P_{a}\lambda}{\hbar c}.$ (8)
This attenuated pulse is detected by the SPD. The time of this detection is registered to produce photon arrival time-based random numbers.
[…]
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Fig. 3. In Fig. 3(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate random numbers as follows.
$\begin{cases}\text{The $i$-th bit is taken as $0$, if }\Delta t_{2i-1}<\Delta t_{2i},\\ \text{The $i$-th bit is taken as $1$, if }\Delta t_{2i-1}>\Delta t_{2i}.\end{cases}$ (9)
But in this process Dynes et al. (2008), the random number generation rate is very slow. The length of the random bit-string is half of the photon count. Comparatively faster processes have been developed in the literature Stipčević and Rogina (2007); Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Wahl et al. (2011); Nie et al. (2014, 2014). In Fig. 3(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by
$P(t_{w}>t)$ $=P(\text{no photon in }(0,t])$
$=P(0,t)=e^{-\mu t}.$ (10)
Therefore, the random number generated with this procedure follows exponential distribution Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Wahl et al. (2011); Nie et al. (2014). Another QRNG scheme is proposed in Stipčević and Rogina (2007), as depicted in Fig. 3(III), using external time reference. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
$=\frac{1}{N^{n}}.$ (11)
V Our Proposal of Two-fold Statistical Method for Testing Quantumness